OpenStudy (zappy620):

Need help with understanding derivative for ((kx^3)^1/2

OpenStudy (zappy620):

\[\sqrt{kx^3}\]

OpenStudy (zappy620):

So i need help with finding the derivative of that.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{d}{dx} \sqrt{x}=\frac{d}{dx}x^{1/2}=\frac{1}{2}x^{(1/2)-1}=\frac{1}{2\displaystyle\sqrt{x}}}\)

OpenStudy (zappy620):

Um actually no. According to Khan Academy is \[\frac{ 3 }{ 2}\sqrt{kx}\]

OpenStudy (zappy620):

Unless that wasnt your final answer :o

OpenStudy (zappy620):

oh nvm i get what your saying

OpenStudy (solomonzelman):

If that answer is correct, I am pretty sure I can come up with this answer too:)

OpenStudy (solomonzelman):

You do know, as I said that \(\color{black}{\displaystyle \frac{d}{dx} \sqrt{x}=\frac{1}{2\displaystyle\sqrt{x}}}\) and thus, by the chain rule principle, in general you have: \(\color{black}{\displaystyle \frac{d}{dx} \sqrt{f(x)}=\frac{1}{2\displaystyle\sqrt{f(x)}}\times f'(x)=\frac{f'(x)}{2\displaystyle\sqrt{f(x)}}}\) and thus, in your particular case you have: \(\color{black}{\displaystyle \frac{d}{dx} \sqrt{kx^3}=\frac{(kx^3)'}{2\displaystyle\sqrt{kx^3}}}\)

OpenStudy (solomonzelman):

and this (if you differentiate correctly) does simplify to the result that you have according to khans academy.

OpenStudy (solomonzelman):

(assuming \(x\in\mathbb{R}\) of course, because \(i^2\ne\displaystyle \sqrt{i^2}\) .)

OpenStudy (zappy620):

Well this particular question was a "Power Rule Challenge" and it was before I would start learning about the chain rule. I'm just confused as to what do when I reach this point. \[\frac{ d }{ dx }(k ^{1/2}x ^{3/2})\]

OpenStudy (zappy620):

I could show all my steps before if you would like as well

OpenStudy (solomonzelman):

Yes, indeed \(\color{black}{\displaystyle \sqrt{kx^3}=\displaystyle k^{1/2}x^{3/2}}\).

OpenStudy (solomonzelman):

So, if you don't want to use the chain rule, you are also fine with that. \(\color{black}{\displaystyle \frac{d}{dx} \sqrt{kx^3}=\displaystyle \frac{d}{dx}k^{1/2}x^{3/2}=\left(k^{1/2}\right)\frac{d}{dx}x^{3/2}}\)

OpenStudy (zappy620):

Yes

OpenStudy (solomonzelman):

But, avoiding the chain rule is not a way out, because there are countless problems that you will very likely have to do for which Chain Rule Would be the ONLY way.

OpenStudy (zappy620):

Yes I understand that. I'm actually moving on to learning that but this particular question has been bothering me.

OpenStudy (solomonzelman):

but, now, this question is clear, isn't it?

OpenStudy (solomonzelman):

Well, for the very least, you know how to differentiate \(k^{1/2}x^{3/2}\). Right?

OpenStudy (zappy620):

So the next step it would be... \[k ^{1/2}\frac{ d }{ dx }(\frac{ 3 }{ 2 }x ^{1/2})\]

OpenStudy (zappy620):

oh i see my error looking at it now :o

OpenStudy (solomonzelman):

you don't need `d/dx` once you have already differentiated. and yes, \(\color{black}{\displaystyle \frac{d}{dx}(k^{1/2}x^{3/2})=\sqrt{k}\times (3/2)x^{1/2}=\frac{3}{2}\sqrt{kx}}\).

OpenStudy (zappy620):

Yea silly me lol. I see it now :D

OpenStudy (zappy620):

But what happens to the x^1/2?

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle x^{1/2}=\sqrt{x}}\)

OpenStudy (zappy620):

Oh i see. \[\sqrt{k}\sqrt{x}=\sqrt{kx}\]

OpenStudy (solomonzelman):

Yup, and again assuming \(k>0\) and \(x>0\).

OpenStudy (solomonzelman):

Just posting this "assumptions" if you like preciseness. (If not, just ignore them.)

OpenStudy (solomonzelman):

oh I mean \(k,x\ge0\) .

OpenStudy (zappy620):

Haha ok. Thank you so much for clearing this up for me. I really appreciate it :D

OpenStudy (solomonzelman):

anyway, if you were to do it with the chain rule, then \(\color{black}{\displaystyle \frac{d}{dx} \sqrt{kx^3}=\frac{(kx^3)'}{2\displaystyle\sqrt{kx^3}}=\frac{3kx^2}{2\displaystyle\sqrt{kx^3}}=\frac{3kx^2}{2\displaystyle\sqrt{kx^2}\times \sqrt{x}}=\frac{3kx^2}{2\displaystyle\sqrt{{\tiny~~}k{\tiny~}}\color{blue}{x}\times \sqrt{x}}.}\) This \(x\) in BLUE, would be \(\sqrt{x^2}=|x|\), but since \(x\ge0\) anyway, because you have an \(x^3\) inside the square root which would demand such a condition. Then, \(\color{black}{\displaystyle \frac{3kx^2}{2\displaystyle\sqrt{{\tiny~~}k{\tiny~}}x \sqrt{x}}=\frac{3}{2}\times\frac{k}{\sqrt{k}}\times\frac{x^2}{\sqrt{x \sqrt{x}}}=\frac{3}{2}\sqrt{k}\times \sqrt{x}=\frac{3}{2}\sqrt{kx} }\)

OpenStudy (zappy620):

Alright good to know. I'll be sure to look back at this after I learn the chain rule in a few :D

OpenStudy (solomonzelman):

Good Luck!